In a country for the year 2019, the minimum consumption level is Rs. 200 and the MPC is .80, the investment is Rs. 250 crores while Government expenditure is Rs. 200 crores. The Government is imposing tax at the rate of 10% of income. Find the equilibrium level of income in the economy and also the amount of consumption, saving and taxes.


1
Expert's answer
2021-12-02T08:59:00-0500

Solution:

At equilibrium level of income: Y = AD

Y = C + I + G

Tax = 0.1Y

C = 200 + 0.8Yd = 200 + 0.8(Y – T) = 200 + 0.8(Y – 0.1Y)

C = 200 + 0.8(Y – 0.1Y)

Y = 200 + 0.8(Y – 0.1Y) + 250 + 200

Y = 200 + 0.8Y – 0.08Y + 250 + 200

Y = 200 + 250 + 200 + 0.8Y – 0.08Y

Y = 650 + 0.72Y

Y – 0.72Y = 650

0.28Y = 650

Y = 2,321.43

The equilibrium level of income (Y) = 2,321

 

Consumption = 200 + 0.8(Y – 0.1Y) = 200 + 0.8(2,321 – 0.1(2,321))

 = 200 + 1856.8 – 232.1 = 1,824.70

Consumption = 1,824.70

Savings = Income - Consumption

Savings = 2,321 – 1824.70 = 496.30

 

Taxes = 0.1Y = 0.1(2,321) = 232.1